An angle is 33 degrees less than one-half its supplement. Find the angle.
Answer by TimothyLamb(4379)

x = angle measure in degrees
---
(180 - x)/2 - 33 = x
90 - x/2 - 33 = x
x + x/2 = 90 - 33
2x/2 + x/2 = 57
3x/2 = 57
x = 57*2/3
x = 38
---
Free algebra tutoring live chat:
https://sooeet.com/chat.php?gn=algebra
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations with quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php
---